OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points.P(6, 2), Q(8, -4)

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

Find the slope between the two using: \[slope = \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

Then plug into this formula: \[y-y_1=m(x-x_1)\] (m is the slope)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You can plug either point into the formula, both will work.